# Planning from Imagination: Episodic Simulation and Episodic Memory for Vision-and-Language Navigation

### v1

@@ 3.2 Recurrent Imagination Tree
We put forth a recurrent local imagination tree mechanism. Drawing inspiration from (Koh et al. 2021; Mallya et al. 2020), the SALI agent can generate high-resolution future image information. For navigation memory $G_{t}$ at step $t$, a length-$K$ history information $H_{t}=\{d_{t-K:t},s_{t-K:t},p_{t-K:t}\}$, the RGB $r_{t}$, and the neighboring position of the current point $p^{g}_{t}$ are utilized to initialize imagery tree $T_{t}^{0}=\{H_{t},r_{t},p^{g}_{t}\}$. The imagery tree $T_{t}^{M}$ is then generated iteratively and finally integrated with the $G_{t}$ as mentioned in Section 3.1.1.
@@ 3.2.1 Recurrent Imagination Mechanism
The generation of the imagination tree is an iterative process. Without loss of generality, Fig 3 illustrates the process of expanding a recurrent imagination tree of $t+N+1$ ($N+1\leq M$) from $t+N$. We first use the input queue $T^{N}_{t}$ to generate a list of structured label images $s_{t+N+1}$ and $d_{t+N+1}$ at moment $t+N+1$ by the inpaint model for each navigable position $p^{g}_{t+N}$. Subsequently, the guidance image $g_{t+N}$ is generated through a point cloud projector. Finally, the RGB image $r_{t+N+1}$ and the target position $p^{g}_{t+N+1}$ will be generated by the spade model with $g_{t+N}$, $s_{t+N+1}$, and $d_{t+N+1}$ as inputs. The short-term memory is then updated to the state $T^{N+1}_{t}$. It is noteworthy that all images are uniformly oriented with 00 headings to prevent the generation of different images for the same point due to variations in agent action orientations. And we set the history length $K$ as 2 for the imagination part.
@@ 3.2.2 High-fidelity Image Generation
1) Inpaint Model. The inpaint model is an encoder-decoder-based network designed for structured label image generation. According to different $p^{g}_{t+N}$, the initial step will convert input data, represented by $H_{t}$, into corresponding structural maps $p_{t+N+1}$ through a point cloud projector. Subsequently, the map is fed into an encoder-decoder based on RedNet and ResNet (He et al. 2016) to generate one-hot semantic code $o_{t+N+1}$ and $d_{t+N+1}$. The structure of the inpaint model is as follows.
@@
$[d_{t+1},o_{t+1}]$ $=\sigma C_{S}(X)+(1-\sigma)C_{F}(X_{t})$ (6)
$\sigma$ is denoted as weight parameter $C_{G}(X_{t})$. $C_{G},C_{S},C_{F}$ stand for Gate-, Shortcut- and Final-Convolution layer.
@@
Nevertheless, the items in the generated semantic image will become fuzzy and illusory with higher $M$ (as illustrated in Fig 4) (Kaneko and Harada 2021). The likelihood of different items in the same pixel would be similar. For this reason, we incorporate the room type model. Given the high correlation between room information and items (e.g., refrigerator-kitchen, bed-bedroom), the SALI agent can ascertain the presence of specific items by determining the room type based on commonsense knowledge. We get the object weight dictionary $w$ for each room and revise the $o_{t+1}$ as $o_{t+1}\cdot(1+w)$. Then it’s transformed to get $s_{t+1}$.
2) Spade Model. The spade model is an image-to-image translation GAN (Wang et al. 2018a; Goodfellow et al. 2020) that enables high-resolution RGB generation capability (Park et al. 2019). The input to the spade model includes the embeddings of $r_{t+N}$ and $g_{t+N}$, and the output is the RGB image $r_{t+N+1}$.
For the following waypoint $p^{g}_{t+N+1}$ prediction, learned $r_{t+N+1}$ and $d_{t+N+1}$ are fed into a waypoint model, a prediction model based on bidirectional BERT.
@@ 3.2.3 Cross-correction
The mechanisms of the global memory and local imagination are mutually supportive. For the global memory, the imagination can assist in enhancing the image features of unvisited nodes ( [Uncaptioned image] ) (Section 3.1.1) and expanding the network structure. For the local imagination, the global map provides the necessary historical information and multiple trajectories, reducing information loss, for a certain imagination point.
@@ 3.3.1 Multimodal Transformer Pre-training
Previous studies have shown the benefits of initializing VLN models with auxiliary tasks. Accordingly, we pre-train our model using expert behavior and imitation learning. The tasks are as follows: masked language modeling (MLM), masked region classification (MRC) (Lu et al. 2019), single-step action prediction (SAP) (Chen et al. 2021), and object grounding (OG) (Lin, Li, and Yu 2021).
@@ 3.3.2 Imagination Model pre-training
1) Inpaint Model. We trained the inpaint model to minimize losses of a cross-entropy loss and a mean absolute error (MAE) term. During the training process, random noise is added to the input.
$\mathcal{L}_{inpaint}=-\lambda\sum\nolimits_{i}\hat{s}_{i}\log(s_{t})+(1-\lambda)\|d_{t}-\hat{d}_{t}\|_{1}$ (7)
$\lambda$ is weight parameter. $s_{t},d_{t}$ are generated semantic- and depth-images, while $\hat{s}_{t},\hat{d}_{t}$ are ground-truth images.
For room-type model training, we get the room-type one-hot codes for each viewpoint from Matterport3D. Two ResNet-50 networks (Russakovsky et al. 2015), which have been previously trained on the ImageNet dataset and VLN data, are employed to encode the semantic image $s$ and the depth image $d$, respectively. The room-type model takes in the encoded feature and outputs one-hot codes. The loss is the cross-entropy between the predicted label $l$ and true room type label $\hat{l}$.
$\mathcal{L}_{room}=-\lambda\sum\nolimits_{i}\hat{l}_{i}\log(l_{i})$ (8)
2) Spade Model. The spade model is a GAN-based model. For the generator, we employ GAN hinge loss, feature matching loss, and perceptual loss (Johnson, Alahi, and Fei-Fei 2016). For PatchGAN-based discriminator (Phillip et al. 2017), the loss function is the loss for both ground-truth images $\hat{r}_{t}$ and generated images $r_{t}$. $\phi^{i}$ and $D^{i}$ denote the output of the $i$-th layer of the pre-trained VGG-19 model and discriminator.
@@
$+\lambda_{P}\sum\nolimits_{i}^{n}\frac{\|D^{i}(\hat{r}_{t})-D^{i}(r_{t})\|_{1}}{n}$ (10)
The waypoint model is a BERT-based waypoint predictor (Hong et al. 2022). For each viewpoint, we labeled its neighboring points’ relative headings and distances into a 120 $\times$ 12 matrix representing 360 degrees and 3 meters (each element represents 3 degrees and 0.25 meters). A heat map, designated as $\hat{m}\in\mathbb{R}^{120\times 12}$, is then generated as the ground-truth data through interpolation of the aforementioned matrix. For the training process, the above-mentioned two ResNet-50 networks are employed to encode the RGB image and depth image for embedding, $e^{r}$ and $e^{d}$. Subsequently, $e^{r},e^{d}$ are merged by a non-linear layer and fed into the model, which generates a heatmap, $m\in\mathbb{R}^{120\times 12}$. Subsequently, non-maximum suppression (NMS) (Hosang, Benenson, and Schiele 2017) is applied to obtain neighboring waypoints. The loss is as follows.
$\mathcal{L}_{waypint}=\frac{1}{120\times 12}\sum\nolimits_{i=1}^{120}\sum\nolimits_{j=1}^{12}(m_{ij}-\hat{m}_{ij})^{2}$ (11)
@@ 3.3.3 Inference
At each stage of the testing process, the online agent constructs a global map of an imagination-based memory. It then reasons about the next action over the map, as explained in Section 3.1. As the memory expands, the SALI agent will approach the correct point and find the shortest path to the target. The agent will stop if it selects the ‘stop’ node or reaches the number limit of action steps. The best model checkpoint is selected by SPL$+$SR.
@@ 4.1 Task Setup
Datasets. We evaluate SALI on VLN benchmarks with both fine-grained instructions (R2R) (Anderson et al. 2018) and high-level ones (REVERIE) (Qi et al. 2020). R2R provides step-by-step instructions with an average length of 32 words. REVERIE gives pre-defined object bounding boxes and instructions to guide agents in describing target objects’ positions. Instructions consist of 21 words on average.
Evaluation Metrics. We adopt the evaluation metrics used commonly by existing works (Qi et al. 2020): 1) Navigation Error (NE) calculates the distance between stop locations and target ones; 2) Trajectory Length (TL) represents the average path length; 3) Success Rate (SR) shows the ratio of stopping within 3m to the target;4) Success rate weighted by Path Length (SPL) makes trade-off between SR and TL; 5) Oracle Success Rate (OSR) is the ratio of including a viewpoint along the path where the target position can be seen; 6) Remote Grounding Success (RGS) is the ratio of successfully executed instructions; 7) RGSPL is RGS penalized by path length.
@@ 4.2 Comparison with State-of-the-art
R2R. Table 1 compares the navigation performance of the SALI agent with a range of approaches on R2R. The SALI agent outperforms the other approaches in all matrices. It is notable that SR increases concurrently with SPL, indicating that the SALI facilitates more correlation between scenes and instruction, thereby enhancing navigation efficiency.
REVERIE. SALI also achieves the best performance on the REVERIE benchmark. SALI surpasses the previous best model. We also note that the RGS and RGSPL increase a lot. We attribute it to the fact that the imagination model, especially the room-type model can help the agent recognize objects around itself better.
@@ 4.3 Quantitative and Qualitative Analysis
In navigation, the agent’s capacity for episodic memory and episodic simulation is reflected in its grasp of the environment-related commonsense. This is particularly evident in the agent’s ability to process complex associative information (e.g., room and object type associations, location associations between rooms).
Quantitative Study. We examine agents’ navigation ability under instructions with complex item and room information. We divided the R2R val_unseen instruction set into:
  • $\mathcal{S}_{1}$: Instructions have over 2 room words and only room words. (e.g. “Walk through hallway and towards restroom.” )
  • $\mathcal{S}_{2}$: Instructions have over 2 object words and only object words (e.g. “Turn left at the oven and past the fridge.” )
  • $\mathcal{S}_{3}$: Complex instructions have over 4 object and room words. (e.g. “Past the door next to the TV in the living room, then walk into the kitchen.” )
We then tested the SALI agent on the three instruction sets as illustrated in Fig 5. Compared to the previous best space-aware model (BEVBert, (An et al. 2023)), SALI wins the most (bandwidth represents a performance gap) on $\mathcal{S}_{3}$ in SR and SPL. SALI’s capabilities of imaging spatial information enable itself to effectively comprehend the interrelationship between objects and rooms, thus performing better when instruction comprises intricate item and room information. We demonstrated SALI achieves commonsense ability through our imagination-memory mechanism, which is analogous to the human episodic simulation memory.
@@
Qualitative Analysis. We visualize the process of generating spatial and image information by the imagination mechanism in Fig 6. The left side shows the imagination for different distance goal positions and the corresponding real images. The peak signal-to-noise ratio (PSNR) (Korhonen and You 2012) of the generated image and the Pearson correlation coefficient (Cohen et al. 2009) with the real image were calculated. The curved area in the figure reflects the mean and variance of the pixels’ error of the two images, which can be approximated by visualizing the Pearson correlation coefficient and PSNR. For spatial waypoint prediction, the predicted heat map $m_{t}$ after the nms operation (top right) is visualized to show neighboring navigable positions.
@@ 4.4 Ablation Study
We conduct extensive experiments on the R2R val-unseen split to evaluate our key design choices.
Imagination vs. Reality. We first investigated the role of the imagined-memory hybrid mechanism in improving navigation performance, see Table 2. It indicates that the memory-based imagery mechanism gets the SoTA performance, which is 12% and 9% higher than “reality only” in SR and SPL. At the same time, we find that navigation using only imagination is the least effective. We attribute this to the inability of the agent to effectively use and correlate the imagination results with the navigation experience and trajectory. This suggests it’s necessary to establish long-term memory mechanisms supporting imaginative abilities.
@@
Temporal and Spatial Imagery. We investigated the effect of the spatio-temporal range of memory on navigation ability. We use the navigation output time step $M$ and the upper limit of the number of nodes $\bar{N}$ to represent the memory range, see Table 3 (Input length is fixed: $K=2$). We can tell that the navigation ability increases with the expansion of the memory. If the navigation range is too large ($M=2,\bar{N}=8$), however, the SPL will decrease. This is because too large a range makes agents at different positions imagine the same destination, which interferes with the agent to make the optimal decision. Furthermore, it is important to consider the trade-off between navigation performance and training time per epoch. However, the overall inference time per epoch is relatively consistent and does not significantly impact the results.
@@
Auxiliary Model Effectiveness. The room type model and the waypoint model in the imagination model are auxiliary models that are not necessary steps for image generation. Waypoint predictions can be made by setting random directions instead. We examined the effect of the two models, see Table 4. We found the agent containing the auxiliary model is optimal in the SPL metrics, reflecting that the two models enable efficient decision-making.
@@
Dynamic Decisioning. The effect of imagination and real node embeddings on the navigation decision is weighted. Comparative experiments are designed, one with dynamic weights (Section 3.2.2) and the other with fixed weights. We observed that dynamic weighting is more effective, with the imagination module’s weight decreasing as navigation progresses, which suggests the later stages of navigation rely less on imagination, akin to human memory mechanism.

### v2

@@ 3.2 Recurrent Imagination Tree
We put forth a recurrent imagination module with a tree structure. Drawing inspiration from (Koh et al. 2021; Mallya et al. 2020), SALI can generate high-resolution future images. At step $t$, a length-$K$ history information $H_{t}=\{d_{t-K:t},s_{t-K:t},p_{t-K:t}\}$, the RGB image $r_{t}$, and the neighboring position of the current position $p^{g}_{t}$—all extracted from navigation memory $G_{t}$—are utilized to initialize imaginary tree $T_{t}^{0}=\{H_{t},r_{t},p^{g}_{t}\}$. The imaginary tree $T_{t}^{M}$ is then generated iteratively and finally integrated with the $G_{t}$ as mentioned in Section 3.1.1.
@@ 3.2.1 Recurrent Imagination Mechanism
The imagination tree grows iteratively. Without loss of generality, Figure 3 illustrates the process of expanding a recurrent imagination tree of imagination step $N+1$ ($N+1\leq M$) from step $t+N$. We first use the input queue $T^{N}_{t}$ to generate a list of structured label images $s^{N+1}_{t}$ and $d_{t}^{N+1}$ at imagination step $N+1$ by the inpaint model for each navigable position $p^{g,N}_{t}$. Subsequently, the guidance image $g^{N}_{t}$ is generated through a point cloud projector. Finally, the RGB image $r_{t}^{N+1}$ and the target position $p^{g,N+1}_{t}$ will be generated by the spade model with the guidance image $g_{t}^{N}$, the semantic image $s_{t}^{N+1}$, and the depth image $d_{t}^{N+1}$ as inputs. The short-term memory is then updated to the state $T^{N+1}_{t}$. It is noteworthy that all images are uniformly oriented with 00 headings to prevent the generation of different images for the same point due to variations in agent action orientations.
@@ 3.2.2 High-fidelity Image Generation
1) Inpaint Model. The inpaint model is an encoder-decoder-based network designed for depth and semantic image generation. For each navigable position $p^{g,N}_{t}$, the process begins by converting historical data $H_{t}$ into corresponding structural maps $x_{t}^{N+1}$ through a point cloud projector. Subsequently, the map is fed into an encoder-decoder based on RedNet and ResNet (He et al. 2016a) to generate one-hot semantic code $e_{t}^{N+1}$ and $d_{t}^{N+1}$, obtained by:
@@
$[d_{t+1},e_{t+1}]$ $=\sigma C_{S}(X)+(1-\sigma)C_{F}(X_{t}),$ (6)
where $\sigma$ is denoted as weight parameter $C_{G}(X_{t})$. $C_{G},C_{S},C_{F}$ stand for Gate-, Shortcut- and Final-Convolution layer.
@@
As the imagination step upper bound $M$ increases, the generated semantic images become increasingly ambiguous, with overlapping likelihoods for objects in the same pixel (Figure 4) (Kaneko and Harada 2021). The likelihood of different objects in the same pixel would be similar. For this reason, we incorporate the room type model. Given the high correlation between room information and objects (e.g., refrigerator-kitchen, bed-bedroom), SALI can ascertain the presence of specific objects by determining the room type based on commonsense knowledge. SALI uses a predefined object weight dictionary $w$ for each room, refining one-hot semantic code $e_{t+1}$ as $e_{t+1}\cdot(1+w)$. The updated code is then transformed to generate $s_{t+1}$.
2) Spade Model. The spade model is an image-to-image translation GAN (Wang et al. 2018a; Goodfellow et al. 2020) that enables high-resolution RGB generation capability (Park et al. 2019). The input to the spade model includes the embeddings of the RGB image $r_{t}^{N}$ and the guidance image $g_{t}^{N}$, and the output is the RGB image $r_{t}^{N+1}$.
For the following navigable position prediction $p_{t}^{g,N+1}$, the generated RGB-D images $r_{t}^{N+1}$ and $d_{t}^{N+1}$ are fed into a BERT-based (Kenton and Toutanova 2019) waypoint model.
@@ 3.2.3 Cross-correction
Global memory and local imagination mechanisms are designed to complement each other. Imagination enhances the image features of navigable nodes and facilitates map expansion. Conversely, the global map provides historical context and trajectory options, reducing information loss for imagination processes.
@@ 3.3.1 Multimodal Transformer Pre-training
Previous studies have shown the benefits of pre-training VLN models with auxiliary tasks. We pre-train our model using expert behavior and imitation learning. Pre-training tasks include masked language modeling (MLM) (Kenton and Toutanova 2019), masked region classification (MRC) (Lu et al. 2019), single-step action prediction (SAP) (Chen et al. 2021), and object grounding (OG) (Lin, Li, and Yu 2021).
@@ 3.3.2 Imagination Model Pre-training
1) Inpaint Model. With random noise added to the input, we trained the inpaint model by minimizing a combination of cross-entropy loss and mean absolute error (MAE):
$\mathcal{L}_{inpaint}=-\lambda\sum\nolimits_{i}\hat{s}_{i}\log(s_{t})+(1-\lambda)\|d_{t}-\hat{d}_{t}\|_{1},$ (7)
where $\lambda$ is weight parameter. $s_{t},d_{t}$ are generated semantic and depth images, while $\hat{s}_{t},\hat{d}_{t}$ are ground-truth images.
For room-type model training, we get the one-hot codes of room-type for each viewpoint from the Matterport3D Simulator (Anderson et al. 2018b). Two ResNet-50 networks (He et al. 2016a), which have been previously trained on the ImageNet dataset (Russakovsky et al. 2015), are employed to encode the semantic image $s$ and the depth image $d$, respectively. The room-type model takes in the encoded feature and outputs one-hot codes. We calculate cross-entropy loss between the predicted label $l$ and true room type label $\hat{l}$:
$\mathcal{L}_{room}=-\lambda\sum\nolimits_{i}\hat{l}_{i}\log(l_{i}).$ (8)
2) Spade Model. The spade model is a GAN-based model. We train the generator with GAN hinge loss, feature matching loss, and perceptual loss (Koh et al. 2021). For PatchGAN-based discriminator (Phillip et al. 2017), we calculate the loss for ground-truth images $\hat{r}_{t}$ and generated images $r_{t}$, where $\phi^{i}$ and $D^{i}$ denote the output of the $i$-th layer of the pre-trained VGG-19 model and discriminator:
@@
$+\lambda_{P}\sum\nolimits_{i}^{n}\frac{\|D^{i}(\hat{r}_{t})-D^{i}(r_{t})\|_{1}}{n}.$ (10)
We implement the waypoint prediction model following (Hong et al. 2022). For each viewpoint, we labeled its neighboring points’ relative headings and distances into a 120 $\times$ 12 matrix representing 360 degrees and 3 meters (each element represents 3 degrees and 0.25 meters). A heat map, designated as $\hat{m}\in\mathbb{R}^{120\times 12}$, is then generated as the ground-truth data through interpolation of the aforementioned matrix. For the training process, the above-mentioned two ResNet-50 networks are employed to encode the RGB image and depth image for embedding, $e^{r}$ and $e^{d}$. Subsequently, $e^{r},e^{d}$ are merged by a non-linear layer and fed into the model, which generates a heatmap, $m\in\mathbb{R}^{120\times 12}$. Subsequently, non-maximum suppression (NMS) (Hosang, Benenson, and Schiele 2017) is applied to obtain neighboring waypoints. The loss is calculated by:
$\mathcal{L}_{waypint}=\frac{1}{120\times 12}\sum\nolimits_{i=1}^{120}\sum\nolimits_{j=1}^{12}(m_{ij}-\hat{m}_{ij})^{2}.$ (11)
@@ 3.3.3 Inference
During inference, the agent constructs a global map of an imagination-based memory on-the-fly. It then reasons about the next action over the map, as explained in Section 3.1. As the memory expands, SALI will approach the correct point and find the shortest path to the target. The agent will stop if it selects the ‘stop’ node or reaches the max action steps.
@@ 4.1 Task Setup
Datasets. We evaluate SALI on VLN benchmarks with both fine-grained instructions (R2R) (Anderson et al. 2018b) and coarse ones (REVERIE) (Qi et al. 2020). R2R provides step-by-step instructions with an average length of 32 words. REVERIE gives pre-defined object bounding boxes and instructions to guide agents in describing target objects’ positions. Instructions consist of 21 words on average.
Evaluation Metrics. We adopt the evaluation metrics (Anderson et al. 2018a) used commonly by existing works: 1) Navigation Error (NE) calculates the distance between stop locations and target ones; 2) Trajectory Length (TL) represents the average path length; 3) Success Rate (SR) shows the ratio of stopping within 3m to the target; 4) Success rate weighted by Path Length (SPL) makes trade-off between SR and TL; 5) Oracle Success Rate (OSR) is the ratio of including a viewpoint along the path where the target position can be seen; 6) Remote Grounding Success (RGS) is the ratio of successfully executed instructions; 7) RGSPL is RGS penalized by path length.
[…]
@@ 4.2 Comparison with State-of-the-Art
R2R. Table 1 compares SALI’s navigation performance with various state-of-the-art approaches on the R2R benchmark. SALI achieves superior performance across all metrics. Notably, the concurrent improvement in SR and SPL demonstrates that SALI effectively aligns scenes with instructions, resulting in more efficient navigation.
REVERIE. SALI also outperforms all previous models on the REVERIE benchmark, achieving the highest scores for RGS and RGSPL. This improvement is attributed to the imagination module, which enhances the agent’s ability to recognize and associate objects within its environment.
@@ 4.3 Quantitative and Qualitative Analysis
SALI’s episodic memory and simulation capabilities are evidenced by its ability to process complex environment-related commonsense, including associations between rooms, objects, and their spatial relationships.
Quantitative Study. We examine agents’ navigation performance under instructions with complex object and room information. We divided the R2R validation unseen instruction set into:
  • $\mathcal{S}_{1}$: Instructions with over two room-related terms and no object terms (e.g. “Walk through hallway and towards restroom.” ).
  • $\mathcal{S}_{2}$: Instructions with over two object-related terms and no room terms (e.g. “Turn left at the oven and past the fridge.” ).
  • $\mathcal{S}_{3}$: Complex instructions with over four combined room and object terms (e.g. “Past the door next to the TV in the living room, then walk into the kitchen.” )
Performance on these instruction sets is illustrated in Figure 5. Compared to the prior best space-aware model (BEVBert, (An et al. 2023)), SALI achieves the largest performance improvement on $\mathcal{S}_{3}$ particularly in SR and SPL. SALI’s ability to imagine spatial relationships allows it to effectively interpret intricate object-room associations, demonstrating improved navigation under complex instructions. This showcases SALI’s commonsense reasoning capabilities, achieved through its imagination-memory mechanism, analogous to human episodic simulation memory.
@@
Qualitative Analysis. Figure 6 visualizes the imagination mechanism’s process of generating spatial and image information. On the left, the imagined images for goal positions at varying distances are compared with corresponding real images. The peak signal-to-noise ratio (PSNR) (Korhonen and You 2012) of the generated image and the Pearson correlation coefficient (Cohen et al. 2009) with the real image were calculated. The shaded curve region represents the mean and variance of pixel errors of the two images, which can be approximated by visualizing the Pearson correlation coefficient and PSNR. For spatial waypoint prediction, the predicted heat map $m_{t}$ after the non-maximum suppression (NMS) operation (top right) is visualized to show neighboring navigable positions.
@@ 4.4 Ablation Study
Extensive experiments on the R2R val-unseen split were conducted to evaluate the proposed features.
Imagination vs. Reality. We first investigated the role of the hybrid mechanism between imagination and memory in improving navigation performance, as shown in Table 2. It indicates that the memory-based imaginary mechanism gets the SoTA performance, which is 12% and 9% higher than “reality only” in SR and SPL. At the same time, we find that navigation using only imagination is the least effective. We attribute this to the inability of the agent to effectively use and correlate the imagination results with its experience and trajectory. This suggests it’s necessary to establish long-term memory mechanisms supporting imaginative abilities.
@@
Temporal and Spatial Imaginary. We investigated the impact of the spatial-temporal memory range on navigation performance, where the navigation output time step $M$ and the upper node limit $\bar{N}$ define the memory range (Table 3), fixed input length $K=2$. Results show that performance improves as the memory range expands. However, excessively large ranges ($M=2,\bar{N}=8$) lead to a decline in SPL, as agents at different positions tend to imagine the same destination, interfering with optimal decision-making. While navigation performance must be balanced with training time per epoch, inference time remains stable and does not significantly affect the results.
@@
Auxiliary Model Effectiveness. The room type model and the waypoint model in the imagination model are auxiliary models that are not necessary for image generation. Waypoint predictions can be made by setting random directions instead. We examined the effect of the two models, as shown in Table 4. We found the agent containing the auxiliary model is optimal in the SPL metrics, reflecting that these two models are essential for decision-making.
@@
Dynamic Decision-making. We evaluate the impact of combining imagined and real node embeddings on navigation performance using static and dynamic weighting strategies. Results show that dynamic weighting is more effective, as the weight of the imagination module decreases over time, reflecting reduced reliance on imagination in later navigation stages, analogous to human memory mechanisms.